OpenStudy (anonymous):

Compute the following binomial probabilities directly from the formula for b(x; n, p): a. b(3; 8, .35) b. b(5; 8, .6) c. P(3<=X<=5) when n=7 and p = .6 d. P(1 <= X) when n=9 and p= .1

OpenStudy (amistre64):

8c3 p^3 q^5

OpenStudy (amistre64):

p = .35 q = 1-p = .65

OpenStudy (amistre64):

for b. its the same process 8c5 p^5 q^3

OpenStudy (amistre64):

c is to find the value at 5 and subtract the value from 3

OpenStudy (amistre64):

or rather; 3+4+5

OpenStudy (amistre64):

i was confusing cumulative with point :)

OpenStudy (amistre64):

does any of this ring a bell?

OpenStudy (anonymous):

Thanks for replying, amistre. I don't understand how to use the formula :(

OpenStudy (amistre64):

the formula deals with factorial/combonatoric notation

OpenStudy (amistre64):

we find the binomial coefficient which tells us how many times the particular set up can occur. 8! 8.7.6.5.4.3.2 8.7.6 ----- = ------------- = ---- = 8.7 = 56 3!5! 3.2 5.4.3.2 3.2 the even occurs 56 times; now we have to just determine the probability of a single event

OpenStudy (amistre64):

a single event that has the set up of 3 successes (p) and the rest failures (q) is the product of p.p.p and q.q.q.q.q

OpenStudy (amistre64):

since p = .35; q is the complement of it; .65 .35*.35*.35* .65*.65*.65*.65*.65* (.35)^3 * (.65)^5

OpenStudy (amistre64):

all together we get: 56 * (.35)^3 * (.65)^5

OpenStudy (amistre64):

that is the basis for the formula; now if any of it doesnt make sense, just say so

OpenStudy (amistre64):

and just to chk: 56 * (.35)^3 * (.65)^5 = .27858... and the table from the wolf gives us: 3 = .2786

OpenStudy (anonymous):

ok, I get it now. The examples in my textbook were throwing me off, according to which I was going to do (8 choose 3) * (.35)^3 * (.35)^3

OpenStudy (amistre64):

that would have been close :) that last bit tho is "whats left over" from the success. in other words: \[\Large ^8C_3\ * p^3\ * (1-p)^{8-3}\]

OpenStudy (amistre64):

b(5; 8, .6); p=.6, so q=.4 \[\Large ^8C_5\ * p^5\ * q^{8-5}\]

OpenStudy (anonymous):

Gives you same answer as part a?

OpenStudy (amistre64):

P(3<=X<=5) when n=7 and p = .6; q=.4 \[\Large ^7C_3\ * p^3\ * q^{7-3}\] \[\Large +^7C_4\ * p^4\ * q^{7-4}\] \[\Large +^7C_5\ * p^5\ * q^{7-5}\]

OpenStudy (amistre64):

... urghh. let me try that again but with the actual problem instead

OpenStudy (amistre64):

56 * .6^5 * .4^3 = .2787 so yeah :)

OpenStudy (amistre64):

the process is the same; it just so happens that the answers are a close fit by coincidence

OpenStudy (anonymous):

awesome! and I understand part c from what you did above. Thanks so much for explaining!

OpenStudy (amistre64):

P(1 <= X) when n=9 and p= .1; q=.9 \[\Large ^9C_0\ * p^0\ * q^{9-0}\]\[\Large +^9C_1\ * p^1\ * q^{9-1}\] youre welcome :)

OpenStudy (amistre64):

do you see my error in the last part here?

OpenStudy (amistre64):

i read X as less than or equal to 1; and summed up the probabilities for the complement instead; which is less work and fixable

OpenStudy (amistre64):

if we are looking for a value of say: A and we find its complement: A'; then since A + A' = 1 we can find A A = 1 - A'

OpenStudy (amistre64):

let me know if that makes sense :)

OpenStudy (anonymous):

so would my answer for part d be approximately .7748?

OpenStudy (anonymous):

ok, it makes sense. Thanks again amistre!!